OpenStudy (anonymous):

Convert to fraction. 15.4%

OpenStudy (anonymous):

15.4 = 15 + 2/5, which is 77/5 Since it is a percent (per cent, means per 100), you have to divide by 100.. so 77/500

OpenStudy (mertsj):

15.4%=15.4/100=154/1000=77/500